OpenStudy (anonymous):

Find the absolute maximum and absolute minimum values of f(x)= e^x -x on interval [-1,1]

OpenStudy (anonymous):

f(x)= e^x -x take the derivative: e^x-1 set equal to zero 0 = e^x-1 solve for x

OpenStudy (anonymous):

thats how far i got, i cannot figure out what to do when i get e^x=1

OpenStudy (anonymous):

this one is a trick of knowing zero powers, any number raised to a power of zero = 1 for example 45^0 = 1 1^0 = 1 makes sense?

OpenStudy (anonymous):

yes so is it just one?

OpenStudy (anonymous):

wow, openstudy is messed up today, sorry i didn't see your respsonse

OpenStudy (anonymous):

e^x=1 so, e^0=1, because any number raised to a zero power =1, so, x=0